# Merging Rate of Opinions via Optimal Transport on Random Measures

### v2

@@ 4.2 Cox completely random measures
Let $\tilde{\mu}$ be a Cox CRM such that $\tilde{\mu}|\tilde{\nu}\sim\textup{CRM}(\tilde{\nu})$ has a.s. finite mean and infinite activity. Then we define its Cox scaled CRM $\tilde{\mu}_{\mathcal{S}}$ as
@@
Following Proposition 3, we see that if $\tilde{\mu}|\tilde{\nu}\sim\mathrm{CRM}(\tilde{\nu})$, then it is a Cox scaled CRM if $\tilde{\nu}\in\mathcal{M}_{1}((0,+\infty)\times\mathbb{X})$ a.s. We denote by $\tilde{\nu}_{\mathcal{S}}$ the random Lévy measure of the Cox scaled CRM, also referred to as random scaled Lévy measure.
@@
Theorem 10 is the analogue for Cox CRMs of Corollary 8 for CRMs. However, it is worth noting that the latter does not follow directly from Theorem 10. Indeed, when dealing with CRMs we do not need the extra assumption that $P_{0}$ is not purely atomic.
By reasoning as for CRMs, thanks to Theorem 10 when comparing normalized Cox CRMs we should restrict their scaled version. Theorem 6 ensures that $\mathcal{W}_{\mathrm{AD}}(\mathcal{L}(\tilde{\nu}^{1}_{\mathcal{S}}),\mathcal{L}(\tilde{\nu}^{2}_{\mathcal{S}}))$ is a distance on random scaled Lévy measures that satisfies $\mathcal{W}_{\mathrm{BL}}(\mathcal{L}(\tilde{\mu}^{1}_{\mathcal{S}}),\mathcal{L}(\tilde{\mu}^{2}_{\mathcal{S}}))\leq\mathcal{W}_{\mathrm{AD}}(\mathcal{L}(\tilde{\nu}^{1}_{\mathcal{S}}),\mathcal{L}(\tilde{\nu}^{2}_{\mathcal{S}}))$.
@@ 5 Merging rate of Bayesian nonparametric models
The remarkable Theorem 1 in James et al. (2009) provides a regular version of the posterior $\mathcal{L}(\tilde{\mu}|X_{1:n})$ arising from model (13), that is, a Markov kernel $\phi:\mathbb{X}^{n}\times\mathcal{B}(\mathcal{M})\to[0,1]$, where $\mathcal{B}(\mathcal{M})$ denotes the Borel $\sigma$-algebra on the space of locally finite measures $\mathcal{M}(\mathbb{X})$. For observed data $x_{1:n}$, we denote by $\mathcal{L}(\tilde{\mu}|X_{1:n}=x_{1:n})$ the probability on $\mathcal{M}(\mathbb{X})$ defined by $A\mapsto\phi(x_{1:n},A)$. A crucial aspect of our setting is that $x_{1:n}$ is a generic deterministic sequence. It could be a realization from the model, a realization from another model, or indeed any deterministic sequence of data: we are able to analyze the asymptotic behaviour of the merging without putting any distributional assumptions on $x_{1:n}$. We will explain how our results depend only on the number of observations $n$ and the number $k=k_{n}$ of distinct values among the $n$ first observations.
We denote by $\tilde{\mu}^{*}\sim\mathcal{L}(\tilde{\mu}|X_{1:n}=x_{1:n})$ a random measure that we often address as the posterior random measure, omitting the dependence on $n$ for notational convenience. The aforementioned Theorem 1 in James et al. (2009), which we state below for completeness, shows that $\tilde{\mu}^{*}$ is a CRM conditionally on a latent variable. Using the terminology of Section 3, this implies that $\tilde{\mu}^{*}$ is a Cox CRM. To preserve the identifiability discussed in Section 4, we do not define the distance at the level of the posterior Cox CRM $\tilde{\mu}^{*}$ but rather on its scaled version $\tilde{\mu}^{i*}_{\mathcal{S}}$ (cf. Definition 7). We focus on the distance $\mathcal{W}_{\mathrm{AD}}$ between their random Lévy measures, where $\mathcal{W}_{\mathrm{AD}}$ is the Wasserstein distance over the $\mathrm{AD}$ metric in Definition 4, which is tailored for Cox scaled CRMs and whose explicit expressions will allow us to derive exact asymptotic behaviours. As described in (12), these can be used to derive upper bounds of $\mathcal{W}_{\mathrm{BL}}$, the Wasserstein distance over the $\mathrm{BL}$ metric in (5), which is a natural distance for arbitrary random measures without additional structure.
@@ 5.1 Dirichlet process
In this section we study the impact of the parameters of the Dirichlet process, which is a normalized gamma CRM, by studying the OT distance between scaled posteriors. We first show that the posterior is a Cox CRM and find its corresponding scaled version (Lemma 12). Then, we prove several finite-sample and asymptotic properties for the OT distance (Proposition 14). This allows us to state new conclusions about the merging rate of opinions (Corollary 15) and study its finite-sample behaviour.
First of all we show that the posterior is a Cox CRM. This result builds on Theorem 11, by further showing that the jumps $J_{i}^{U}$ are infinitely divisible, and can thus be included in the random Lévy intensity, as shown in Lemma SM4 in the supplementary material.
@@
Lemma 12 shows that the posterior of a gamma CRM is a Cox CRM with a peculiar feature: all the randomness of its Lévy intensity is contained in a scale parameter, so that the corresponding Cox scaled CRM reduces to a scaled CRM. In the sequel we write $\Gamma(z,t)=\int_{t}^{+\infty}x^{1-z}e^{-x}\,\mathrm{d}x$ for the incomplete Gamma function.
@@
Let $\tilde{\mu}^{i}$ be a gamma CRM of parameters $(\alpha_{i},P_{0}^{i})$ with $P_{0}^{i}$ having finite first moment for $i=1,2$. Then, $\mathcal{W}_{\mathrm{AD}}(\mathcal{L}(\tilde{\nu}^{1}_{\mathcal{S}}),\mathcal{L}(\tilde{\nu}^{2}_{\mathcal{S}}))=\mathrm{AD}(\nu^{1*}_{\mathcal{S}},\nu^{2*}_{\mathcal{S}})=\mathcal{J}+\mathcal{A}$, with $\mathcal{J}=\mathcal{J}(\alpha_{1},\alpha_{2},n)$ and $\mathcal{A}=\mathcal{A}(\alpha_{1},P_{0}^{1},\alpha_{2},P_{0}^{2},n)$ defined as
$\mathcal{J}$ $=\int_{0}^{+\infty}|(\alpha_{1}+n)\Gamma(0,(\alpha_{1}+n)s)-(\alpha_{2}+n)\Gamma(0,(\alpha_{2}+n)s)|\mathrm{d}s;$
@@
Let $\tilde{\mu}^{1},\tilde{\mu}^{1}$ be two gamma CRMs of parameters $(\alpha_{i},P_{0}^{i})$, for $i=1,2$, with $\alpha_{1}\neq\alpha_{2}$ or $P_{0}^{1}\neq P_{0}^{2}$. Then for any sequence $(x_{n})_{n\geq 1}$ of observations on $\mathbb{X}$,
@@
Theorem 15 states that there will always be merging of opinions, and provides us with the merging rate of $1/n$, which holds independently of the choice of parameters $(\alpha_{i},P_{0}^{i})$ and independently of the dimension of the space $\mathbb{X}$. Moreover, thanks to Proposition 14 we can also study the starting time of the merging, which highlights a different behaviour for the jump and the atom component, as discussed in the remainder of this section.
Figure 1: Jump component $\mathcal{J}$ of $\mathcal{W}_{\mathrm{AD}}(\mathcal{L}(\tilde{\nu}^{1*}_{\mathcal{S}}),\mathcal{L}(\tilde{\nu}^{2*}_{\mathcal{S}}))$ when $\tilde{\mu}^{1}$ is a gamma$(\alpha_{1}=1,P_{0}^{1})$ CRM and $\tilde{\mu}^{2}$ is a gamma$(\alpha_{2},P_{0}^{2})$ CRM, as the number of observations $n$ increases. The upper bounds are the right hand side of (19).
Point 1. of Proposition 14 states that the jump component $\mathcal{J}$ decreases steadily in $n$, that is, the merging starts with the first observation (time 1). We illustrate this behavior in Figure 1, where without loss of generality we fix $\alpha=1$ and plot the value of $\mathcal{J}$ as a function of $n$, for different values of $\alpha_{2}$, together with the upper bound (19), which becomes asymptotically exact as $n\to+\infty$, as stated in point 3. of Proposition 14. As for the atom component $\mathcal{A}$, this can show different behaviours depending on the parameters $(\alpha_{i},P_{0}^{i})$, as showcased by the study of the upper bounds in point 4. of Proposition 14. In particular, if $\alpha_{1}=\alpha_{2}=\alpha$,
@@
Thus in this scenario $\mathcal{A}$ steadily decreases as $n$ increases, that is, the merging of the atoms starts at time 1. The left panel of Figure 2 validates it numerically by taking $\alpha_{1}=\alpha_{2}=\alpha$ for different values of $\alpha$, $P_{0}^{1}$ and $P_{0}^{2}$ Gaussian distributions with different mean and same variance, and i.i.d. observations from a Poisson distribution.
Figure 2: Atom component $\mathcal{A}$ of $\mathcal{W}_{\mathrm{AD}}(\mathcal{L}(\tilde{\nu}^{1*}_{\mathcal{S}}),\mathcal{L}(\tilde{\nu}^{2*}_{\mathcal{S}}))$, as the number of observations $n$ increases, in two different scenarios. Left: $\tilde{\mu}^{1}$ is a gamma$(\alpha_{1}=\alpha,P_{0}^{1}=\mathcal{N}(1,1))$ CRM and $\tilde{\mu}^{2}$ is a gamma$(\alpha_{2}=\alpha,P_{0}^{2}=\mathcal{N}(2,1))$ CRM, where $\mathcal{N}(m,v)$ is a Gaussian distribution of mean $m$ and variance $v$. Right: $\tilde{\mu}^{1}$ is a gamma$(\alpha_{1}=10,P_{0}^{1}=\mathcal{N}(1,1))$ CRM and $\tilde{\mu}^{2}$ is a gamma$(\alpha_{2},P_{0}^{2}=\mathcal{N}(1,1))$ CRM. In both scenarios the observations are an i.i.d. sample from a Poisson distribution of mean 1.
@@
For many standard data-generating mechanisms for the data, i.e., if we consider $(x_{n})_{n\geq 1}$ as a realization of a random sequence $(X_{n})_{n\geq 1}$ possibly differing from model (13), $\mathcal{W}_{d_{\mathbb{X},\text{t}}}(P_{0},\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}})$ has a non-zero limit almost surely. For example, this happens if the data are i.i.d. samples from $P\neq P_{0}$ thanks to the law of large numbers or if they are exchangeable thanks to de Finetti Theorem. In such case, for $n$ large enough, the behaviour of $\mathcal{A}$ with respect to $n$ reduces to the behaviour of the prefactor
@@
This quantity can be studied analytically: it first increases and reaches a maximum at $n\approx\sqrt{\alpha_{1}\alpha_{2}}$ before decreasing to 00 at rate $1/n$, hinting that this may well be the case for $\mathcal{A}$ as well, at least when $\sqrt{\alpha_{1}\alpha_{2}}$ is large enough. We tested this behavior numerically. In Figure 2 on the right we fix $P_{0}^{1}=P_{0}^{2}=P_{0}$ to be a Gaussian distribution, $\alpha_{1}=10$ and $\alpha_{2}$ to vary, while the observations are i.i.d. from $P$ a Poisson distribution: in this setting we know that the expected value of $\mathcal{W}_{d_{\mathbb{X},\text{t}}}(P_{0},\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}})$ converges to $\mathcal{W}_{d_{\mathbb{X},\text{t}}}(P_{0},P)$ with rate $1/\sqrt{n}$ (Fournier and Guillin, 2015). We see that the distance is not decreasing in $n$, the maximum being achieved for $\bar{n}\approx 34$ when $\alpha_{2}=50$, $\bar{n}\approx 42$ when $\alpha_{2}=100$ and $\bar{n}\approx 64$ when $\alpha_{2}=500$. As expected, the maximum is better approximated by $\sqrt{\alpha_{1}\alpha_{2}}$ (equal to approximately 21, 31, and 72, respectively) as $\sqrt{\alpha_{1}\alpha_{2}}$ increases.
In conclusion, depending on the total base measure and the base probabilities, we can have different behaviours: either a steady decrease, or an increase phase followed by a decrease, but ultimately the asymptotic rate of convergence is $1/n$. From a statistical perspective this finding is quite remarkable: even if two Bayesians have the same prior on a phenomenon, the difference between their opinions may increase if they see a finite amount of data, according to how much confidence they are willing to put on their prior guesses.

### v3

@@ 4.2 Cox completely random measures
[…]
@@
[…]
@@
Theorem 10 is the analog for Cox CRMs of Corollary 8 for CRMs. However, it is worth noting that the latter does not follow directly from Theorem 10. Indeed, when dealing with CRMs we do not need the extra assumption that $P_{0}$ is not purely atomic.
By reasoning as for CRMs, thanks to Theorem 10 when comparing normalized Cox CRMs we should restrict our attention to their scaled version. Theorem 6 ensures that $\mathcal{W}_{\mathrm{AD}}(\mathcal{L}(\tilde{\nu}^{1}_{\mathcal{S}}),\mathcal{L}(\tilde{\nu}^{2}_{\mathcal{S}}))$ is a distance on random scaled Lévy measures satisfying $\mathcal{W}_{\mathrm{BL}}(\mathcal{L}(\tilde{\mu}^{1}_{\mathcal{S}}),\mathcal{L}(\tilde{\mu}^{2}_{\mathcal{S}}))\leq\mathcal{W}_{\mathrm{AD}}(\mathcal{L}(\tilde{\nu}^{1}_{\mathcal{S}}),\mathcal{L}(\tilde{\nu}^{2}_{\mathcal{S}}))$.
@@ 5 Merging rate of Bayesian nonparametric models
The remarkable Theorem 1 in James et al. (2009) provides a regular version of the posterior $\mathcal{L}(\tilde{\mu}|X_{1:n})$ arising from model (13), that is, a Markov kernel $\phi:\mathbb{X}^{n}\times\mathcal{B}(\mathcal{M})\to[0,1]$, where $\mathcal{B}(\mathcal{M})$ denotes the Borel $\sigma$-algebra on the space of locally finite measures $\mathcal{M}(\mathbb{X})$. For observed data $x_{1:n}$, we denote by $\mathcal{L}(\tilde{\mu}|X_{1:n}=x_{1:n})$ the probability measure on $\mathcal{M}(\mathbb{X})$ defined by $A\mapsto\phi(x_{1:n},A)$. A crucial aspect of our setting is that $x_{1:n}$ is a generic deterministic sequence. It could be a realization from the model, a realization from another model, or indeed any deterministic sequence of data: we are able to analyze the asymptotic behavior of the merging without putting any distributional assumptions on $x_{1:n}$. We will explain how our results depend only on the number of observations $n$ and the number $k=k_{n}$ of distinct values among the $n$ first observations.
We denote by $\tilde{\mu}^{*}\sim\mathcal{L}(\tilde{\mu}|X_{1:n}=x_{1:n})$ a random measure that we often address as the posterior random measure, omitting the dependence on $n$ for notational convenience. The aforementioned Theorem 1 in James et al. (2009), which we state below for completeness, shows that $\tilde{\mu}^{*}$ is a CRM conditionally on a latent variable. Using the terminology of Section 3, this implies that $\tilde{\mu}^{*}$ is a Cox CRM. To preserve the identifiability discussed in Section 4, we do not define the distance at the level of the posterior Cox CRM $\tilde{\mu}^{*}$ but rather on its scaled version $\tilde{\mu}^{*}_{\mathcal{S}}$ (cf. Definition 7). We focus on the distance $\mathcal{W}_{\mathrm{AD}}$ between their random Lévy measures, where $\mathcal{W}_{\mathrm{AD}}$ is the Wasserstein distance over the $\mathrm{AD}$ metric in Definition 4, which is tailored for Cox scaled CRMs and whose explicit expressions will allow us to derive exact asymptotic behaviors. As described in (12), these can be used to derive upper bounds of $\mathcal{W}_{\mathrm{BL}}$, the Wasserstein distance over the $\mathrm{BL}$ metric in (5), which is a natural distance for arbitrary random measures without additional structure.
@@ 5.1 Dirichlet process
In this section we study the impact of the parameters of the Dirichlet process, which is a normalized gamma CRM, by studying the OT distance between scaled posteriors. We first show that the posterior is a Cox CRM and find its corresponding scaled version (Lemma 12). Then, we prove several finite-sample and asymptotic properties for the OT distance (Proposition 14). This allows us to state new conclusions about the merging rate of opinions (Corollary 15) and study its finite-sample behavior.
First of all we show that the posterior is a Cox CRM. This result builds on Theorem 11, by further showing that the jumps $J_{i}^{U}$ are infinitely divisible, and can thus be included in the random Lévy intensity, as shown in Lemma SM4 in the Supplementary Material (Catalano and Lavenant, 2025).
@@
Lemma 12 shows that the posterior of a gamma CRM is a Cox CRM with a peculiar feature: all the randomness of its Lévy intensity is contained in a scale parameter, so that the corresponding Cox scaled CRM reduces to a scaled CRM. In the sequel we write $\Gamma(z,t)=\int_{t}^{+\infty}x^{z-1}e^{-x}\,\mathrm{d}x$ for the incomplete Gamma function.
@@
Let $\tilde{\mu}^{i}$ be a gamma CRM of parameters $(\alpha_{i},P_{0}^{i})$ for $i=1,2$. Then it holds that $\mathcal{W}_{\mathrm{AD}}(\mathcal{L}(\tilde{\nu}^{1}_{\mathcal{S}}),\mathcal{L}(\tilde{\nu}^{2}_{\mathcal{S}}))=\mathrm{AD}(\nu^{1*}_{\mathcal{S}},\nu^{2*}_{\mathcal{S}})=\mathcal{J}+\mathcal{A}$, with $\mathcal{J}=\mathcal{J}(\alpha_{1},\alpha_{2},n)$ and $\mathcal{A}=\mathcal{A}(\alpha_{1},P_{0}^{1},\alpha_{2},P_{0}^{2},n)$ defined as
$\mathcal{J}$ $=\int_{0}^{+\infty}|(\alpha_{1}+n)\Gamma(0,(\alpha_{1}+n)s)-(\alpha_{2}+n)\Gamma(0,(\alpha_{2}+n)s)|\mathrm{d}s,$
@@
Let $\tilde{\mu}^{1},\tilde{\mu}^{2}$ be two gamma CRMs of parameters $(\alpha_{i},P_{0}^{i})$, for $i=1,2$, with $\alpha_{1}\neq\alpha_{2}$ or $P_{0}^{1}\neq P_{0}^{2}$. Then for any sequence $(x_{n})_{n\geq 1}$ of observations on $\mathbb{X}$,
@@
Theorem 15 states that there will always be merging of opinions, and provides us with the merging rate of $1/n$, which holds independently of the choice of parameters $(\alpha_{i},P_{0}^{i})$ and independently of the dimension of the space $\mathbb{X}$. Tracking down the multiplicative constants in the proof of Theorem 15, we actually prove the upper bound
$\mathrm{AD}(\nu^{1*}_{\mathcal{S}},\nu^{2*}_{\mathcal{S}})\leq\frac{1}{n}\left\{\min(\alpha_{1},\alpha_{2})\mathcal{W}_{d_{\mathbb{X},\text{t}}}(P_{0}^{1},P_{0}^{2})+(c+2)|\alpha_{2}-\alpha_{1}|\right\}.$
Thus, even if the difference in the priors (measured by $\mathcal{W}_{d_{\mathbb{X},\text{t}}}(P_{0}^{1},P_{0}^{2})$ and $|\alpha_{2}-\alpha_{1}|$) does not affect the rate, it influences the premultiplicative factor. Moreover, thanks to Proposition 14 we can also study the starting time of the merging, which highlights a different behavior for the jump and the atom component, as discussed in the remainder of this section.
[…]
Point 1. of Proposition 14 states that the jump component $\mathcal{J}$ decreases steadily in $n$, that is, the merging starts with the first observation (time 1). We illustrate this behavior in Figure 1, where without loss of generality we fix $\alpha=1$ and plot the value of $\mathcal{J}$ as a function of $n$, for different values of $\alpha_{2}$, together with the upper bound (19), which becomes asymptotically exact as $n\to+\infty$, as stated in point 3. of Proposition 14. As for the atom component $\mathcal{A}$, this can show different behaviors depending on the parameters $(\alpha_{i},P_{0}^{i})$, as showcased by the study of the upper bounds in point 4. of Proposition 14. In particular, if $\alpha_{1}=\alpha_{2}=\alpha$,
@@
Thus, in this scenario $\mathcal{A}$ steadily decreases as $n$ increases, that is, the merging of the atoms starts at time 1. The left panel of Figure 2 validates it numerically by taking $\alpha_{1}=\alpha_{2}=\alpha$ for different values of $\alpha$, $P_{0}^{1}$ and $P_{0}^{2}$ Gaussian distributions with different mean and same variance, and i.i.d. observations from a Poisson distribution.
[…]
@@
For many standard data-generating mechanisms for the data, i.e., if we consider $(x_{n})_{n\geq 1}$ as a realization of a random sequence $(X_{n})_{n\geq 1}$ possibly differing from model (13), $\mathcal{W}_{d_{\mathbb{X},\text{t}}}(P_{0},\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}})$ has a non-zero limit almost surely. For example, this happens if the data are i.i.d. samples from $P\neq P_{0}$ thanks to the law of large numbers or if they are exchangeable thanks to de Finetti Theorem. In such case, for $n$ large enough, the behavior of $\mathcal{A}$ with respect to $n$ reduces to the behavior of the prefactor
@@
This quantity can be studied analytically: it first increases and reaches a maximum at $n\approx\sqrt{\alpha_{1}\alpha_{2}}$ before decreasing to 0 at rate $1/n$, hinting that this may well be the case for $\mathcal{A}$ as well, at least when $\sqrt{\alpha_{1}\alpha_{2}}$ is large enough. We tested this behavior numerically. In Figure 2 on the right we fix $P_{0}^{1}=P_{0}^{2}=P_{0}$ to be a Gaussian distribution, $\alpha_{1}=10$ and $\alpha_{2}$ to vary, while the observations are i.i.d. from $P$ a Poisson distribution: in this setting we know that the expected value of $\mathcal{W}_{d_{\mathbb{X},\text{t}}}(P_{0},\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}})$ converges to $\mathcal{W}_{d_{\mathbb{X},\text{t}}}(P_{0},P)$ with rate $1/\sqrt{n}$ (Fournier and Guillin, 2015). We see that the distance is not decreasing in $n$, the maximum being achieved for $\bar{n}\approx 34$ when $\alpha_{2}=50$, $\bar{n}\approx 42$ when $\alpha_{2}=100$ and $\bar{n}\approx 64$ when $\alpha_{2}=500$. As expected, the maximum is better approximated by $\sqrt{\alpha_{1}\alpha_{2}}$ (equal to approximately 21, 31, and 72, respectively) as $\sqrt{\alpha_{1}\alpha_{2}}$ increases.
The previous paragraph shows that if $P^{1}_{0}=P^{2}_{0}$ but $\alpha_{1}\neq\alpha_{2}$ then $\mathcal{A}$ may be increasing. The quantity $\mathcal{A}$ corresponds to the Wasserstein distance between the predictive distributions $\mathcal{L}(X_{n+1}|X_{1:n})$ if $(X_{n})_{n\geq 1}$ follow model (1), thus is directly relevant to the practitioner. Moreover, we can also guarantee that for some values of the parameters and some data the distance $\mathrm{AD}(\nu^{1*}_{\mathcal{S}},\nu^{2*}_{\mathcal{S}})$ must first be increasing. Indeed, as $\mathcal{A}=0$ when $n=0$ and $P^{1}_{0}=P^{2}_{0}$, then $\mathrm{AD}(\nu^{1*}_{\mathcal{S}},\nu^{2*}_{\mathcal{S}})>\mathrm{AD}(\nu^{1}_{\mathcal{S}},\nu^{2}_{\mathcal{S}})$ as soon as
$\mathcal{W}_{d_{\mathbb{X},\text{t}}}\bigg{(}P_{0},\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}}\bigg{)}\geq\frac{(\alpha_{1}+n)(\alpha_{2}+n)}{(\alpha_{2}-\alpha_{1})n}(\mathcal{J}(\alpha 1,\alpha 2,0)-\mathcal{J}(\alpha 1,\alpha 2,n)),$
meaning that the Lévy intensities a posteriori are further apart than a priori. For example, if $\alpha_{1}=10$, $\alpha_{2}=50$, and $n=20$, it is the case if $\mathcal{W}_{d_{\mathbb{X},\text{t}}}(P_{0},\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}})\geq 1.012$. However, with our analytical tools we cannot conclude if $\mathcal{W}_{\mathrm{BL}}(\mathcal{L}(\tilde{\mu}^{1*}_{\mathcal{S}}),\mathcal{L}(\tilde{\mu}^{2*}_{\mathcal{S}}))$ can be increasing, being $\mathrm{AD}(\nu^{1*}_{\mathcal{S}},\nu^{2*}_{\mathcal{S}})$ only an upper bound to $\mathcal{W}_{\mathrm{BL}}(\mathcal{L}(\tilde{\mu}^{1*}_{\mathcal{S}}),\mathcal{L}(\tilde{\mu}^{2*}_{\mathcal{S}}))$. Proposition 2 yields the lower bound $\mathcal{W}_{\mathrm{BL}}(\mathcal{L}(\tilde{\mu}^{1*}_{\mathcal{S}}),\mathcal{L}(\tilde{\mu}^{2*}_{\mathcal{S}}))\geq\mathcal{A}/2$, so that both the upper and lower bounds are increasing for small values of $n$. However, this does not necessarily imply that $\mathcal{W}_{\mathrm{BL}}(\mathcal{L}(\tilde{\mu}^{1*}_{\mathcal{S}}),\mathcal{L}(\tilde{\mu}^{2*}_{\mathcal{S}}))$ is also increasing. A formal proof is left for future work.
In conclusion, depending on the total base measure and the base probabilities, we can have different behaviors for the adapted extended Wasserstein discrepancy: either a steady decrease, or an increase phase followed by a decrease, but ultimately the asymptotic rate of convergence is $1/n$. From a statistical perspective this finding is quite remarkable: even if two Bayesians have the same prior on a phenomenon, the difference between their opinions may increase if they see a finite amount of data, according to how much confidence they are willing to put on their prior guesses.